OpenStudy (bowtie_man123):

What is the slope of the line that passes through the points (–5, 3) and (7, 5)? A. –1 B. C. 1 D. 6

OpenStudy (igreen):

Plug the points in the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{5-3}{7+5}\)

OpenStudy (igreen):

Simplify

OpenStudy (anonymous):

I say d